What would be the effect on national saving (S) if there was an equal-sized reduction in...

  1. What would be the effect on national saving (S) if there was an equal-sized reduction in both government purchases (G↓) and personal taxes (T↓)?  Up, Down, or Same?
Answer #1

National savings ( S) = Private sanings + Public savings = ( Y - C - T ) + ( T- G )

S = Y - C - T + T - G = Y - C - G

Equal sized reduction in G and T , so public savings in unchanged ( T - G ) . Private savings rises ( Y - C - T ) .

So we can see that Taxes ( T ) does not affect S . If there is increase or decrease in T , it cancels out .

If G reduces then Savings rises . Answer : Up .